OpenStudy (anonymous):

PLEASEEEEE HELPP! ! ! ! Given: 3x + 1 = 2 + 2x – 4 Prove: x = –3 Given the equation 3x + 1 = 2 + 2x – 4, use the commutative property to rearrange the terms so that like terms are next to one another. This gives the equation 3x + 1 = 2 – 4 + 2x. Then use the associative property of addition to group the like terms.

OpenStudy (anonymous):

. This gives the equation 3x + 1 = 2 – 4 + 2x. Then use the associative property of addition to group the like terms. This gives the equation 3x + 1 = (2 – 4) + 2x. Next, combine like terms to get the equation 3x + 1 = – 2 + 2x. Use the subtraction property of equality to subtract 2x from both sides of the equation. This gives the equation x + 1 = – 2. Then use the _________________________ to subtract 1 from both sides of the equation. This gives the solution x = –3. Therefore, given the equation 3x + 1 = 2 + 2x – 4, x is equal to –3. Which justification was left out of the paragraph proof above?

OpenStudy (mertsj):

Then use the addition property of equalityto add negative 1 to both sides of the equation.

OpenStudy (anonymous):

That or use again the subtraction property of equality and subtract 1 from both sides of the equation. But it's basically the same thing.

OpenStudy (mertsj):

Oh, you mean subtracting is the same as adding the opposite? Who knew.

OpenStudy (anonymous):

Yes; I would say so, Since lets say that you added - 2x to both sides instead of subtracting 2x; you'll get the same thing.

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

you're welcome :)